The Carry It Company makes various size crates. Each crate is in the shape of a topless cube with wooden handles on two sides. IN an effort to save money on building supplies, the company reduces the length of each dimension of a crate by 7%. What is the resulting percent decrease in surface area of each carrying crate/
Answer by LinnW(1048)

If s represents the length of a side of the original crate,
the external surface area of the box is 5 times
the area of squares with length s or
5(s*s).
If the length of a side is reduced by 7%,
the surface area of the new crate is
5((s-7% of s)(s-7% of s)) =
5(0.93s)(0.93s)
= 5(0.8649 s*s)
This means that (1 - 0.8649) = 0.1351 = 13.51 % reduction in surface area.